OpenStudy (anonymous):

Use the ratio test to determine convergence or divergence. Sigma n!/n^n; n = 1; to infinity.

OpenStudy (anonymous):

(xz-zy)^3 evaluate the expression for: x=3,y=2,z=5 (15-10)^3 = 5^3 = 125

OpenStudy (experimentx):

seems to converge

OpenStudy (anonymous):

Yes, it converges, but I have to just show work. Don't know how to start.